OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10? -5/2 5/2 2/5 -2/5 none of the above

OpenStudy (austinl):

You simply solve for y, and then it will be in the form \(y=mx+b\) where \(m=\text{slope}\)

OpenStudy (shamil98):

2x + 5y = -10 5y = -2x - 10 y = -2/5 - 2 slope is -2/5

OpenStudy (shamil98):

BEAT YOU TO IT AUSTIN ,hue.

OpenStudy (austinl):

\(2x+5y=-10\) \(5y=-2x-10\) \(y=-\dfrac{2}{5}x-2\) \(m=-\dfrac{2}{5}\\ b=-2\)

OpenStudy (anonymous):

Damn thanks @shamil98 Damn you good ;)

OpenStudy (shamil98):

you're welcome.